# High-mobility heavy quasiparticles in a van der Waals antiferromagnetic dense Kondo lattice CeTe3

### v1

@@ RESULTS
$\Delta\rho(\mu_{0}H,T)\propto\frac{\lambda T}{\sinh(\lambda T)}e^{-\lambda T_{D}}\cos\left(\frac{2\pi F}{\mu_{0}H}-\pi+\Phi_{B}\right),$ (2)
where $\lambda=(2\pi^{2}k_{B}m^{*})/(\hbar e\mu_{0}H)$, $T_{D}$ is Dingle temperature, and $\Phi_{B}$ is Berry phase. Multiple oscillation frequencies are discernible in Fast Fourier Transform (FFT) spectrum of $\Delta\rho_{xx}(1/\mu_{0}H)$, seeing Figure 3B. Besides the previously reported predominant oscillation frequency $F_{\alpha}=36$ T that shows up at relatively low field ($<8$ T) CeTe3-QO1 ; CeTe3-QO2 ; CeTe3-QO3 , our high-field SdH measurements unveil four additional frequencies, $F_{\beta}=478$ T, $F_{\gamma}=884$ T, $F_{\eta}=1760$ T, and $F_{\delta}=2200$ T. Quantum oscillations were also observed in magnetization measurements, viz dHvA effect, seeing Figure S1C and the inset to Figure 3A. It seems the dHvA effect is more sensitive to those high-frequency oscillations (e.g., $F_{\beta}$), whereas the smallest $F_{\alpha}$ is hardly seen. Table 1 summarizes the information of all the observed quantum oscillations. It is well known that the oscillation frequency is related to the extrema cross-sectional area $A_{F}$ of FS (perpendicular to the applied field) through the Onsager relation $F=(\hbar/2\pi e)A_{F}$ Shoenberg-MagneticOscillations . The observed oscillation frequencies correspond to $0.16\%$, $2.16\%$, $3.99\%$, $7.94\%$ and $9.93\%$ of the entire BZ, respectively. According to previous ARPES studies CeTe3-FS ; NdTe3-FS , $F_{\alpha}$ arises from a small pocket in the butterfly-shaped FS wing near the X point (Figure 1C); $F_{\beta}$ and $F_{\gamma}$ can be ascribed to the elliptical pockets around the X point and are consequences of the Fermi surface reconstruction due to the CDW transition; the largest $F_{\delta}$ likely corresponds to the square pocket in the center of BZ. The exact origin of $F_{\eta}$ is still unclear based on ARPES results NdTe3-FS ; here we tentatively assign it to the section nearby the $\delta$. Another possibility for $F_{\eta}$ is the second harmonics of the $\gamma$ pocket. ARPES with improved resolution will be needed to clarify this issue. It should also be pointed out that the FS structure is slightly modulated across the AFM1 transition, cf. Table 1. We will revisit this issue later.
@@
Sample Pocket Region $F$ (T) $m^{*}~{}(m_{e})$ $\mu_{q}$ (cm2/Vs)
@@
Remarkably, the SdH quantum oscillations of CeTe3 persist to higher than 50 K (Figure 3C), indicative of high mobility of this correlated vdW material. According to the LK formula Eq. (2), the FFT amplitude decays with increasing temperature following a law of $\lambda T/\sinh{\lambda T}$, and the fitting of the experimental data to this formula shall lead to the effective mass $m^{*}$. A major feature here, however, is that the temperature dependence of FFT amplitude exhibits a distinctive inflection around $T_{N1}$, and this is particularly clear for $F_{\alpha}$. Such kind of non-LK temperature dependence has been observed in a handful of materials including mercury ElliottM-JPFMP1980 , fractional quantum Hall systems TsuiD-PRL1982 and SmB6 TanB-Science2015 with different mechanisms. In CeTe3, since the deviation is coincident with the AFM1 transition, it is reasonable to ascribe it to the Fermi surface modulation by magnetic ordering, as mentioned by Higashihara et al CeTe3-QO3 . Separate fittings reveal that the quasiparticle effective mass changes from $m^{*}_{\alpha}=0.035(1)~{}m_{e}$ for above $T_{N1}$ to $m^{*}_{\alpha}=0.46(4)~{}m_{e}$ for below. The Dingle temperature can then be deduced from the LK analysis, $T_{D}=1.90$ K (below $T_{N1}$) and $T_{D}=32.61$ K (above $T_{N1}$), and this in turn yields the quantum lifetime $\tau_{q}=\hbar/2\pi k_{B}T_{D}$ and quantum mobility $\mu_{q}=e\tau_{q}/m^{*}$. It is worthwhile to mention that the enhancement of quasiparticle effective mass below $T_{N1}$ does not cause a reduction of quantum mobility, but rather, $\mu_{q}$ is increased by $27.6\%$, and meanwhile, $\tau_{q}$ is enlarged by about 17 times. Such an enhancement in $\mu_{q}$ and $\tau_{q}$ reminds us that the AFM transition at $T_{N1}$ may be accompanied with a topological transition. For this purpose, a Landau fan plot is constructed for the $\alpha$ FS, and the results of 1.5 K and 5.0 K are compared in Figure 3E. Our high-field SdH measurements allow determining the Berry phase accurately. According to Lifshitz-Onsager quantization rule for 2D system LO-QO : $A_{F}(\hbar/2\pi e\mu_{0}H)=n+1/2-\Phi_{B}/2\pi$, the Berry phase $\Phi_{B}$ can be obtained from the intercept when linearly extrapolating the Lanau level index to the limit of $1/\mu_{0}H\rightarrow 0$. The derived intercepts are 0.047 and $-0.448$ for 1.5 and 5 K, meaning the corresponding Berry phase $\Phi_{B}$ is close to $\pi$ and 0, respectively. The nontrivial Berry phase at 1.5 K suggests that CeTe3 may be a kind of antiferromagnetic topological semimetal due to which the backward scattering rate is greatly suppressed by the topological protection LiangT-Cd3As2 . However, we must admit here that right now it is hard for us to further prove this by first-principles calculations, as the exact magnetic structure below $T_{N1}$ remains unclear to date; ARPES measurements below $T_{N1}$ are also challenging.
@@
(B) Temperature-dependent sheet resistance for D1, D2, and D3 flakes, showing the two sequential AFM transitions at $T_{N1}$ and $T_{N2}$.
@@
(F) Temperature dependence of normalized FFT amplitudes of the $\alpha$ oscillation; comparison among nanoflakes D1, D2 and bulk sample. The black and blue lines represent the LK fits above and below $T_{N1}$, respectively. Note that the experimental data are normalized at 5 K.
@@
(i) The AFM transitions survive even in the thinnest 6L nanoflake, seeing Figure 4b. The critical temperature $T_{N1}$ in D1-3 are only slightly decreased with respect to the bulk, although the signature of the transition in the $R_{s}(T)$ seems much washed-out. To our surprise, the AFM2 transition at $T_{N2}$ is promoted by exfoliation. In bulk samples, $T_{N2}$ was determined by a $\lambda$-shaped peak in specific heat around 1.3 K Magnetic-CeTe3 . The value now shifts to $\sim 1.8$ K in nanoflakes, as manifested by an abrupt discontinuity in $R_{s}(T)$ and a hysteresis loop which are most obvious in D1 (inset to Figure 4B). In other words, the AFM2 transition now becomes 1st-order-like in nature. The mechanism for these modifications remains mysterious for us. One possibility is due to a strong coupling between AFM2 and CDW orders, the latter of which can be promoted by exfoliation as a consequence of FS instability in lower dimensionality GdTe3-Sci adv . Another possibility is that, because the AFM1 transition is softened, as a compensation, the magnetic entropy release at AFM2 is enlarged; the abundance of spin fluctuations near $T_{N2}$ may drive the transition into first-order ManoschekJ-MnSiPRB2013 . Additional experiments such as neutron scattering are required to further clarify this issue.
@@
(ii) The high mobility of the bulk CeTe3 is also inherited in nanoflakes, as indicated by the robust SdH oscillations persisting above 50 K (Figures 4C and S6). The FFT of D1 manifests that all the five oscillation frequencies are observable (Figure 4E) and their values are close to those in the bulk sample, seeing Table 1. D2 and D3 also exhibit pronounced quantum oscillations, but the signals for $\eta$- and $\delta$-pockets are less discernible. To make a good comparison of the effective mass between nanoflake and bulk samples, we normalize their FFT amplitudes of the $\alpha$ pocket at 5 K (Figure 4F). In the paramagnetic phase, all the curves roughly fall onto a single line that is determined by the effective mass $m^{*}=0.036(2)~{}m_{e}$, implying that the extent of mass renormalization caused by Kondo hybridization is nearly the same between bulk and nanoflakes. More strikingly, across $T_{N1}$, the upturn of FFT amplitude in the nanoflakes are more drastic than in the bulk sample, suggesting a stronger band modulation by AFM1 ordering. Indeed, the obtained $m^{*}$ is enhanced from $0.037(1)~{}m_{e}$ (above $T_{N1}$) to $0.70(7)~{}m_{e}$ (below $T_{N1}$), while $\mu_{q}$ is increased by 46.5%; the enhancement ratios of both $m^{*}$ and $\mu_{q}$ are larger than in bulk. This suggests that the scenario of topologically-protected scattering is likely applicable to nanoflakes as well. In particular, $\mu_{q}$ at 1.5 K reaches 3158 cm2/Vs, about 31.4% higher than that of the bulk.
@@ DISCUSSION
Finally, it is necessary to briefly discuss about the classic transport mobility $\mu$ that can be deduced from a two-band model analysis on the Hall conductivity $\sigma_{xy}$ (seeing Figure S4) LuoY-WTe2APL2015 . The obtained temperature dependent carrier concentration ($n_{e,h}$) and classic transport mobility ($\mu_{e,h}$) are shown in Figure S5, where the subscripts $e$ and $h$ are short for electron and hole, respectively. We note that the maximum $\mu_{e}$ reaches approximately 3400 cm2/Vs for bulk sample and 2000 cm2/Vs for D3 nanoflake. This, again, demonstrates the high mobility of CeTe3. An enhancement of $\mu_{e}$ is also observed around $T_{N1}$, resembling that of $\mu_{q}$. The high transport mobility is also reflected by the low scattering rate $1/\tau_{1}$ determined from fitting optical conductivity to Eq. (1), seeing Figure S3. An upturn in $1/\tau_{1}$ is visible below $T_{K}^{on}$, in line with the onset of Kondo scattering revealed by electrical transport.
In conclusion, we present a comprehensive survey of the vdW AFM Kondo lattice CeTe3 using optical spectroscopy and high-field quantum oscillation measurements. Our results not only complete the Fermi surface topology proposed by earlier ARPES experiment, but also confirm the effective mass enhancement by both AFM1-induced band-structure modulation and Kondo-hybridization-induced narrow band. Despite the mass enhancement, the low-temperature quantum mobility surprisingly increases and reaches $\sim$2400-3200 cm2/Vs. The magnetic orders, high mobility and correlation effect are inherited or even promoted in atomically thin nanoflakes. All these findings suggest that vdW CeTe3 is a rare example that simultaneously possesses high mobility and electronic correlation, providing a fascinating platform to study 2D Kondo physics and to manipulate novel quantum states.
@@ METHODS
CeTe3 single crystals were grown using the self-flux method. High-purity raw materials Ce and Te were weighed and mixed in a molar ratio of 3 : 97. The mixture was placed in an alumina crucible, vacuum-sealed in a quartz tube, and heated up to 900 C within 12 hours, held for 12 hours, and then slowly cooled to 550 C at a rate of 2 C/h. After maintaining at 550 C for 6 hours, the excess Te flux was removed by centrifugation. Plate-like single crystals with dimensions up to 3$\times$2$\times$0.5 mm3 were obtained. The obtained single crystals were verified by single-crystal X-ray diffraction using XtaLAB mini II single-crystal diffractometer. This material behaves moderate sensitivity to air and should be stored in an anoxia and dry environment.
@@
The reflectivity $R(\omega)$ of CeTe3 was measured using a Bruker Vertex 80v Fourier transform spectrometer, with the light polarized in the $\mathbf{ac}$ planes. An in-situ gold overfilling technique was used to obtain the absolute $R(\omega)$ at different temperatures between 5 and 300 K over a broad frequency ranging from 50 to 7000 cm-1. $R(\omega)$ from 7000 to 25000 cm-1 was also measured at room temperature. Then, we employed an AvaSpec-2048$\times$14 optical fiber spectrometer to measure $R(\omega)$ in the frequency up to 50000 cm-1 at room temperature. The real part of the optical conductivity $\sigma_{1}(\omega)$ was determined from a Kramers-Kronig analysis of the measured $R(\omega)$. We adopted a Hagen-Rubens ($R=1-A\sqrt{\omega}$) low-frequency extrapolation below 50 cm-1. A constant $R(\omega)$ up to 12.5 eV, followed by a free-electron ($\omega^{-4}$) response, was used for the high-frequency extrapolation.

### v2

@@ RESULTS
$\Delta\rho(\mu_{0}H,T)\propto\frac{\lambda T}{\sinh(\lambda T)}e^{-\lambda T_{\text{D}}}\cos\left(\frac{2\pi F}{\mu_{0}H}-\pi+\Phi_{\text{B}}\right),$ (2)
where $\lambda=(2\pi^{2}k_{B}m^{*})/(\hbar e\mu_{0}H)$, $T_{\text{D}}$ is Dingle temperature, and $\Phi_{\text{B}}$ is Berry phase. Multiple oscillation frequencies are discernible in Fast Fourier Transform (FFT) spectrum of $\Delta\rho_{xx}(1/\mu_{0}H)$, seeing Figure 3B. Besides the previously reported predominant oscillation frequency $F_{\alpha}=36$ T that shows up at relatively low field (<8<8 T) CeTe3-QO1 ; CeTe3-QO2 ; CeTe3-QO3 , our high-field SdH measurements unveil four additional frequencies, $F_{\beta}=478$ T, $F_{\gamma}=884$ T, $F_{\eta}=1760$ T, and $F_{\delta}=2200$ T. Quantum oscillations were also observed in magnetization measurements, viz dHvA effect, seeing Figure S1C and the inset to Figure 3A. It seems the dHvA effect is more sensitive to those high-frequency oscillations (e.g., $F_{\beta}$), whereas the smallest $F_{\alpha}$ is hardly seen. In particular, frequency splitting of $F_{\eta}$ is also distinguishable in dHvA, which is likely a consequence of weak FS warping and bilayer splitting RTe3-breakdown . Table 1 summarizes the information of all the observed quantum oscillations. It is well known that the oscillation frequency is related to the extrema cross-sectional area $A_{F}$ of FS (perpendicular to the applied field) through the Onsager relation $F=(\hbar/2\pi e)A_{F}$ Shoenberg-MagneticOscillations . The observed oscillation frequencies correspond to $0.16\%$, $2.16\%$, $3.99\%$, $7.94\%$ and $9.93\%$ of the entire BZ, respectively. According to previous ARPES studies CeTe3-FS ; NdTe3-FS , $F_{\alpha}$ arises from a small pocket in the butterfly-shaped FS wing near the X point (Figure 1C); $F_{\beta}$ and $F_{\gamma}$ can be ascribed to the elliptical pockets around the X point and are consequences of the Fermi surface reconstruction due to the CDW transition; the largest $F_{\delta}$ likely corresponds to the square pocket in the center of BZ. The exact origin of $F_{\eta}$ is still unclear based on ARPES results NdTe3-FS ; here we tentatively assign it to the section nearby the $\delta$. Another possibility for $F_{\eta}$ is the second harmonics of the $\gamma$ pocket. ARPES with improved resolution will be needed to clarify this issue. It should also be pointed out that the FS structure is slightly modulated across the AFM1 transition, cf. Table 1. We will revisit this issue later.
@@
Sample Pocket Region $F$ (T) $m^{*}~(m_{\text{e}})$ $\mu_{\text{q}}$ (cm2/Vs)
@@
Remarkably, the SdH quantum oscillations of CeTe3 persist to higher than 50 K (Figure 3C), indicative of high mobility of this correlated vdW material. According to the LK formula Eq. (2), the FFT amplitude decays with increasing temperature following a law of $\lambda T/\sinh{\lambda T}$, and the fitting of the experimental data to this formula shall lead to the effective mass $m^{*}$. A major feature here, however, is that the temperature dependence of FFT amplitude exhibits a distinctive inflection around $T_{\text{N1}}$, and this is particularly clear for $F_{\alpha}$ [Figure 3D]. Such kind of non-LK temperature dependence has been observed in a handful of materials including mercury ElliottM-JPFMP1980 , fractional quantum Hall systems TsuiD-PRL1982 and SmB6 TanB-Science2015 with different mechanisms. In CeTe3, since the deviation is coincident with the AFM1 transition, it is reasonable to ascribe it to the Fermi surface modulation by magnetic ordering, as mentioned by Higashihara et al CeTe3-QO3 . Separate fittings reveal that the quasiparticle effective mass changes from $m^{*}_{\alpha}=0.035(1)~m_{\text{e}}$ for above $T_{\text{N1}}$ to $m^{*}_{\alpha}=0.46(4)~m_{\text{e}}$ for below. The Dingle temperature can then be deduced from the LK analysis, $T_{\text{D}}=1.90$ K (below $T_{\text{N1}}$) and $T_{\text{D}}=32.61$ K (above $T_{\text{N1}}$), and this in turn yields the quantum lifetime $\tau_{\text{q}}=\hbar/2\pi k_{\text{B}}T_{\text{D}}$ and quantum mobility $\mu_{\text{q}}=e\tau_{\text{q}}/m^{*}$. It is worthwhile to mention that the enhancement of quasiparticle effective mass below $T_{\text{N1}}$ does not cause a reduction of quantum mobility, but rather, $\mu_{\text{q}}$ is increased by $27.6\%$, and meanwhile, $\tau_{\text{q}}$ is enlarged by about 17 times. Such an enhancement in $\mu_{\text{q}}$ and $\tau_{\text{q}}$ reminds us that the AFM transition at $T_{\text{N1}}$ may be also accompanied with a topological transition. For this purpose, a Landau fan plot is constructed for the $\alpha$ FS, and the results of 1.5 K and 5.0 K are compared in Figure 3E. Our high-field SdH measurements allow determining the Berry phase accurately. For the $\alpha$ pocket, the application of 34 T magnetic field confines the carriers below the $n=2$ Landau level. According to Lifshitz-Onsager quantization rule for 2D system LO-QO : $A_{F}(\hbar/2\pi e\mu_{0}H)=n+1/2-\Phi_{\text{B}}/2\pi$, the Berry phase $\Phi_{\text{B}}$ can be obtained from the intercept when linearly extrapolating the Lanau level index to the limit of $1/\mu_{0}H\rightarrow 0$. The derived intercepts are 0.047 and 0.448-0.448 for 1.5 and 5 K, meaning the corresponding Berry phase $\Phi_{\text{B}}$ is close to $\pi$ and 0, respectively. Additional temperature dependent Berry phase is shown in Figure S8. The nontrivial Berry phase at 1.5 K strongly suggests that CeTe3 may be a kind of antiferromagnetic topological semimetal due to which the backward scattering rate is greatly suppressed by the topological protection LiangT-Cd3As2 . However, we must admit here that right now it is hard for us to further prove this by first-principles calculations, as the exact magnetic structure below $T_{\text{N1}}$ remains unclear to date; ARPES measurements below $T_{\text{N1}}$ are also challenging.
Before closing this section, we shall briefly elaborate the discrepancy in the extent of mass renormalization revealed through quantum oscillation, optical conductivity, and specific heat. Quantum oscillation is FS-dependent and mobility-sensitive, primarily probing the $m^{*}$ of specific light pockets, for instance $m_{\alpha}^{*}$ in our case. Optical conductivity assesses the effective mass of conduction bands through the total Drude weight [cf. the first term of Eq. (1)], meaning that the derived effective mass represents the average of all free charge carriers, denoted as $\langle m^{*}\rangle$. Specific heat, on the other hand, measures the density of states from all the electronic contributions including both conduction and localized electrons. Notably, localized electrons can also enter the specific heat via magnetic entropy (e.g. magnetic transitions, spin fluctuations, Kondo hybridization, Schottky anomalies, and so on) which usually is significant for Ce-$4f$ in moderate Kondo lattice, as is the case in CeTe3 Magnetic-CeTe3 . Although these different techniques yield differing values for effective mass or Sommerfeld coefficient, they consistently point to the presence of considerable electronic correlation effect in CeTe3 at low temperature.
@@
(B) Temperature-dependent sheet resistance for D1, D2, and D3 flakes, showing the two sequential AFM transitions at $T_{\text{N1}}$ and $T_{\text{N2}}$.
@@
(F) Temperature dependence of normalized FFT amplitudes of the $\alpha$ oscillation; comparison among nanoflakes D1, D2 and bulk sample. The black and blue lines represent the LK fits above and below $T_{\text{N1}}$, respectively. Note that the experimental data are normalized at 5 K.
@@
(i) The AFM transitions survive even in the thinnest 6L nanoflake, seeing Figure 4B. The critical temperature $T_{\text{N1}}$ in D1-3 are only slightly decreased with respect to the bulk, although the signature of the transition in the sheet resistance $R_{\text{s}}(T)$ seems much washed-out. To our surprise, the AFM2 transition at $T_{\text{N2}}$ is promoted by exfoliation. In bulk samples, $T_{\text{N2}}$ was determined by a $\lambda$-shaped peak in specific heat around 1.3 K Magnetic-CeTe3 . The value now shifts to $\sim 1.8$ K in nanoflakes, as manifested by an abrupt discontinuity in $R_{\text{s}}(T)$ and a hysteresis loop that is most obvious in D1 (inset to Figure 4B). In other words, the AFM2 transition now becomes 1st-order-like in nature. The mechanism for these modifications remains mysterious for us. One possibility is due to a strong coupling between AFM2 and CDW orders, the latter of which can be promoted by exfoliation as a consequence of FS instability in lower dimensionality GdTe3-Sci adv . Another possibility is that, because the AFM1 transition is softened, as a compensation, the magnetic entropy release at AFM2 is enlarged; the abundance of spin fluctuations near $T_{\text{N2}}$ may drive the transition into first-order ManoschekJ-MnSiPRB2013 . Additional experiments such as neutron scattering are required to further clarify this issue.
@@
(ii) The high mobility of the bulk CeTe3 is also inherited in nanoflakes, as indicated by the robust SdH oscillations persisting above 50 K (Figures 4C and S7). The FFT of D1 manifests that all the five oscillation frequencies are observable (Figure 4E) and their values are close to those in the bulk sample, seeing Table 1. D2 and D3 also exhibit pronounced quantum oscillations, but the signals for $\eta$- and $\delta$-pockets are less discernible. To make a good comparison of the effective mass between nanoflake and bulk samples, we normalize their FFT amplitudes of the $\alpha$ pocket at 5 K (Figure 4F). In the paramagnetic phase, all the curves roughly fall onto a single line that is determined by the effective mass $m^{*}=0.036(2)~m_{\text{e}}$, implying that the extent of mass renormalization caused by Kondo hybridization is nearly the same between bulk and nanoflakes. More strikingly, across $T_{\text{N1}}$, the upturn of FFT amplitude in the nanoflakes are more drastic than in the bulk sample, suggesting a stronger band modulation by AFM1 ordering. Indeed, the obtained $m^{*}$ is enhanced from $0.037(1)~m_{\text{e}}$ (above $T_{\text{N1}}$) to $0.70(7)~m_{\text{e}}$ (below $T_{\text{N1}}$), while $\mu_{\text{q}}$ is increased by 46.5%; the enhancement ratios of both $m^{*}$ and $\mu_{\text{q}}$ are larger than in bulk. This suggests that the scenario of topologically-protected scattering is likely applicable to nanoflakes as well. In particular, $\mu_{\text{q}}$ at 1.5 K reaches 3158 cm2/Vs, about 31.4% higher than that of the bulk.
@@ DISCUSSION
Finally, it is necessary to briefly discuss about the classic transport mobility $\mu$ that can be deduced from a two-band model analysis on the Hall conductivity $\sigma_{xy}$ (seeing Figure S5) LuoY-WTe2APL2015 . The obtained temperature dependent carrier concentration ($n_{\text{e,h}}$) and classic transport mobility ($\mu_{\text{e,h}}$) are shown in Figure S6, where the subscripts e and h are short for electron and hole, respectively. The maximum $\mu_{\text{e}}$ reaches approximately 3400 cm2/Vs for bulk sample and 2000 cm2/Vs for D3 nanoflake at 1.5 K. This, again, demonstrates the high mobility of CeTe3. An enhancement of $\mu_{\text{e}}$ is also observed around $T_{\text{N1}}$, resembling that of $\mu_{\text{q}}$. We note that the values of $\mu_{\text{q}}$ and $\mu_{\text{e}}$ are of the same order of magnitude in CeTe3, hinting that large-angle scattering process is still dominant AndoT-RMP1982 , despite the non-trivial topological nature. This is in sharp contrast to most non-interacting topological materials such as NbAs where $\mu$ is much larger than $\mu_{\text{q}}$ due to the suppression of backward scattering by topological protection LuoY-NbAsSdH . This counter-indicates that Kondo effect plays an important role in the transport properties of CeTe3 GorkovLP-CeIn3PRB2006 ; Haines-CeGeLTP2012 . The high transport mobility in CeTe3 is also reflected by the low scattering rate $1/\tau_{1}$ determined from fitting optical conductivity to Eq. (1), seeing Figure S3. An upturn in $1/\tau_{1}$ is visible below $T_{K}^{on}$, in line with the onset of Kondo scattering revealed by electrical transport.
In conclusion, we present a comprehensive survey of the vdW AFM Kondo lattice CeTe3 using optical spectroscopy and high-field quantum oscillation measurements. Our results not only complete the Fermi surface topology proposed by earlier ARPES experiment, but also confirm the effective mass enhancement by both AFM1-induced band-structure modulation and Kondo-hybridization-induced narrow band. Despite the mass enhancement, the low-temperature quantum mobility surprisingly increases and reaches $\sim$2400-3200 cm2/Vs. The magnetic orders, high mobility and correlation effect are inherited or even promoted in atomically thin nanoflakes. All these findings suggest that vdW CeTe3 is a rare example that possesses both high charge mobility and electronic correlation, providing a fascinating platform to study 2D Kondo physics and to manipulate novel quantum states.
@@ METHODS
[…]
@@
[…]